What is the derivative of "y = arctanx\/ln2x"


1
Expert's answer
2021-06-09T08:19:04-0400

Let us find the derivative of "y =\frac{ \arctan x}{\ln 2x}:"


"y' =\frac{ \frac{1}{1+x^2}\ln 2x-\arctan x\frac{2}{2x}}{\ln^2 2x}=

\frac{x\ln 2x-(1+x^2)\arctan x}{x(1+x^2)\ln^2 2x}."